OpenStudy (anonymous):

HELLPPP! @misty1212 y^2 + 6y - 2x + 13 = 0 Focus: Directrix: Length of focal chord:

OpenStudy (misty1212):

\[y^2 + 6y - 2x + 13 = 0 \]

OpenStudy (anonymous):

That's right

OpenStudy (misty1212):

you know what this looks like?

OpenStudy (misty1212):

you are going to have to do some work to find the vertex, which you need to start the problem first you have to complete the square on the y term

OpenStudy (anonymous):

I found the vertex already, (2, -3)

OpenStudy (misty1212):

oh then you are9/10th of the way done how did you find it?

OpenStudy (anonymous):

A graphing calculator

OpenStudy (misty1212):

oh

OpenStudy (misty1212):

that does not help much

OpenStudy (anonymous):

Why do you need to know HOW I found the vertex?

OpenStudy (misty1212):

\[y^2+6y=2x-13\\ (y+3)^2=2x-13+9\\ (y+3)^2=2x-4\\ (y+3)^2=2(x-2)\] vertex is \((2,-3)\)

OpenStudy (misty1212):

because if you find the vertex by making it look like \[(y+3)^2=2(x-2)\] then finding the rest is easy

OpenStudy (misty1212):

this looks just like \[(y-k)^2=4p(x-h)\] a parabola that opens to the right with vertex \((h,k)\) and what you need is \(p\) in your case \(4p=2\) and so \(p=\frac{1}{2}\) your focus is \(\frac{1}{2}\) to the right of the vertex, the directrix is \(\frac{1}{2}\) units to the left

OpenStudy (anonymous):

What the focal chord?

OpenStudy (misty1212):

what is a focal cord?

OpenStudy (anonymous):

no, what was the focal chord?

OpenStudy (misty1212):

it does not ask for the focal chord, look carefully at the question dear

OpenStudy (misty1212):

there is no one "focal chord" there are infinitely many it asks for the LENGTH of the focal chord that is \(4p\) which in your case is \(2\)

OpenStudy (anonymous):

Ohhhhh, okay. Thanks